OpenStudy (anonymous):

If you put 6 equilateral triangles together as shown below, you can make a regular hexagon. What is the area, in square units, of a regular hexagon of side lenght 6 units?

OpenStudy (anonymous):

hlep me

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

r u still there?

OpenStudy (anonymous):

Find the area of each triangle and add them up

OpenStudy (anonymous):

I know what u mean, helps if y u post a pic.

OpenStudy (anonymous):

It's not my question estudier

OpenStudy (anonymous):

I wasn't talking to u...

OpenStudy (anonymous):

I was just making a suggestion

OpenStudy (anonymous):

Our posts are crossing....

OpenStudy (anonymous):

OK, as Heromiles says, you need to calculate the area of 1 triangle and multiply it by 6. Can u do that?

OpenStudy (anonymous):

She would have to understand how to find the height of the equilateral triangle

OpenStudy (anonymous):

Equilateral means all the sides of the triangle are the same length. If you can find the area of one triangle, you can find the area of the entire hexagon by multiplying the area of one triangle by 6. To find the area of a triangle, you need a base measurement (easy in this case; it is 6) and a height measurement. For equilateral triangles, the height measurement will be the measure of the midpoint of the base of the triangle to the top vertex/point. You can figure out this height using Pythagorean theorem (a^2 = b^2 + c^2). Let me draw a picture, then I'll continue.

OpenStudy (anonymous):

root(3) Area of an Equilateral triangle = ------- * (side)^2 4 Here side of the equilateral triangle will be 6 units root(3) root(3) so Area of 1 Eqltrl trngl = -------- * (6)^2 = -------- * 36 = 9 root(3) 4 4 Since the regular hexagon is made up of such 6 equilateral triangles... its area = 6* 9 root(3) i.e.\[54\sqrt{3}\]

OpenStudy (anonymous):

So, once you've constructed the height, you will notice two right angle triangles. The base measurement for each is 3 units (half the base of the equilateral triangle), and the hypotenuse is 6 cm. Using these numbers in the formula, you get. 6^2 = 3^2 + c^2 36 = 9 + c^2 c^2 = 36 - 9 c^2 = 27 c = sqrt(27) You can now find the area of each triangle using the formula A = 1/2 * bh

OpenStudy (anonymous):

@MathDrills why do we need to go through all this procedure.... we do not need to calculate the height......... since length of side is known, we can use the formula I used...